OpenStudy (anonymous):

estimate the sum of the decimal by rounding to the whole number . 4.493 8.962 4.018+ 4.4=5, 8.3=9,5.2=5

OpenStudy (anonymous):

add them 5+9+5=19 am i right here

OpenStudy (anonymous):

5 + 9 + 4 = 18

OpenStudy (anonymous):

ok what about 16.058- 6.143

OpenStudy (anonymous):

16 - 6 = 10